OpenStudy (anonymous):

Help please The acute angle x radians is such that tanx=k, where k is a positive constant. Express in terms of k... i) tan(pi-x) ii) tan (1/2pi-x) iii)sinx

OpenStudy (anonymous):

2

OpenStudy (anonymous):

tan (pi/2 - x) = tanx

OpenStudy (anonymous):

property

OpenStudy (anonymous):

But the answers 1/k for part two

OpenStudy (anonymous):

oh , u need to evaluate , huh?

OpenStudy (anonymous):

okay...

OpenStudy (anonymous):

1) tan(pi-x) = -tanx = -k 2)

OpenStudy (anonymous):

tan(pi/2 - x) = tan(90degrees - x) = cotx = 1/k

OpenStudy (rina.r):

tan(pi-x) i think is the answer

OpenStudy (anonymous):

sinx / cosx = tanx sinx = tanx * cosx = kcosx

OpenStudy (anonymous):

gottit ?

OpenStudy (anonymous):

Sweet cheers man, appreciate it

OpenStudy (anonymous):

no probs , mate !!